# LiNo: Advancing Recursive Residual Decomposition of Linear and Nonlinear Patterns for Robust Time Series Forecasting

### v1

@@ Abstract
Forecasting models are pivotal in a data-driven world with vast volumes of time series data that appear as a compound of vast Linear and Nonlinear patterns. Recent deep time series forecasting models struggle to utilize seasonal and trend decomposition to separate the entangled components. Such a strategy only explicitly extracts simple linear patterns like trends, leaving the other linear modes and vast unexplored nonlinear patterns to the residual. Their flawed linear and nonlinear feature extraction models and shallow-level decomposition limit their adaptation to the diverse patterns present in real-world scenarios. Given this, we innovate Recursive Residual Decomposition by introducing explicit extraction of both linear and nonlinear patterns. This deeper-level decomposition framework, which is named LiNo, captures linear patterns using a Li block which can be a moving average kernel, and models nonlinear patterns using a No block which can be a Transformer encoder. The extraction of these two patterns is performed alternatively and recursively. To achieve the full potential of LiNo, we develop the current simple linear pattern extractor to a general learnable autoregressive model, and design a novel No block that can handle all essential nonlinear patterns. Remarkably, the proposed LiNo achieves state-of-the-art on thirteen real-world benchmarks under univariate and multivariate forecasting scenarios. Experiments show that current forecasting models can deliver more robust and precise results through this advanced Recursive Residual Decomposition. We hope this work could offer insight into designing more effective forecasting models. Code is available at this Repository: https://github.com/Levi-Ackman/LiNo.
@@ 1 Introduction
Time series forecasting (TSF) is a long-established task (Lim & Zohren, 2021; Wang et al., 2024b), with a wide range of applications (Zhou et al., 2022a; LIU et al., 2022; Piao et al., 2024). Notably, numerous deep learning methods have been employed to address the TSF problem, utilizing architectures such as Recurrent Neural Networks (RNNs) (Elman, 1990; Lin et al., 2023), Temporal Convolutional Networks (TCNs) (donghao & wang xue, 2024; Wu et al., 2023), Multilayer Perceptron (MLP) (Liu et al., 2023; Challu et al., 2023) and Transformers (Liu et al., 2022a; Kitaev et al., 2020). Recent advancements in the time series forecasting community have suggested that seasonal (nonlinear) and trend (linear) decomposition can enhance forecasting performance (Zhang et al., 2022b; Wu et al., 2021). This is typically performed once, with the trend component being extracted using methods such as the simple moving average kernel (MOV) (Zhou et al., 2022b; Wang et al., 2023), the exponential smoothing function (ESF) (Woo et al., 2022), or a learnable 1D convolution kernel (LD) (Yu et al., 2024). The seasonal component is then obtained by subtracting the trend part from the original time series.
@@
Figure 1: Example of the multi-level linear and nonlinear patterns in real-world time series. We take the ETTh2 dataset as an example and decompose a raw time series (Raw) into four signals through linear and nonlinear patterns decomposition. Linear 1&2 is the obtained linear patterns using the proposed Li block, and Nonlinear 1&2 is the obtained nonlinear patterns using No block.
@@
Real-world time series data often exhibit more complex structures, combining multiple levels of linear and nonlinear characteristics (Stock & Watson, 1998; Dama & Sinoquet, 2021; Agrawal & Adhikari, 2013), as illustrated in Figure 1. When deployed in real-world time series data, simple seasonal and trend decomposition faces three main challenges. Firstly, these methods rely on simple techniques (MOV, LD, ESF). MOV and LD employed a fixed window size, which cannot fully extract all linear patterns, while the non-learnability of ESF limits its performance. They can only extract simple linear features such as trends, while other linear features like cyclic patterns and autoregression (Chatfield & Xing, 2019) remain underutilized. Secondly, we argue the obtained seasonal part is actually a residual consisting of all the unextracted linear and nonlinear information. Without further separating the nonlinear part from the residual severely hinders its extraction. Since it’s extremely challenging for deep models to extract useful information from such mixtures (Bengio et al., 2013; Tishby & Zaslavsky, 2015). Another problem is the design of current nonlinear models, which mainly focus on one or two types of nonlinear patterns. Such designs cannot satisfy our requirements in the real world. Thirdly, their shallow-level decomposition is incompatible with the multi-level characteristics of real-world time series. In contrast, numerous studies have shown that deeper-level decomposition can lead to better time series analysis (Huang et al., 1998; Rilling et al., 2003). Therefore, it is crucial to develop advanced decomposition methods capable of multi-granularity separation of various modes, and better linear and nonlinear pattern extractors to provide a more nuanced understanding of the signal’s structure and potentially improve forecasting performance.
@@
To address the challenges above, we propose adopting Recursive Residual Decomposition (RRD), a method used in Empirical Mode Decomposition (EMD) (Huang et al., 1998; Rilling et al., 2003), to decompose a time series into multiple patterns. This process is performed recursively. Each pattern is extracted based on the residual obtained by subtracting previously extracted patterns from the original signal, utilizing Intrinsic Mode Functions (IMF) to identify similar characteristics. Instead of traditional RRD, we explicitly and alternately extract both linear and nonlinear patterns in this paper. Specifically, We can employ a Li block (MOV, LD, ESF, etc.) as the IMF for extracting linear patterns and a No block (Transformer, RNN, etc.) as the IMF for extracting nonlinear patterns. Then, the RRD is performed on a deeper level. We denote the proposed overall framework as LiNo. To fully realize RRD’s potential, we propose an advanced LiNo, Specifically, we enhance the Li block to a general learnable autoregressive model with a full receptive field and propose a novel No block capable of modeling essential nonlinear features, such as temporal variations, period information, and inter-series dependencies in multivariate forecasting.
In summary, our contributions can be delineated as follows.
@@
  • Proposed No block demonstrates better nonlinear pattern extraction ability than current SOTA nonlinear models, such as iTransformer and TSMxier.
  • LiNo consistently delivers top-tier performance in both multivariate and univariate forecasting scenarios, demonstrating robust resilience against various noise disturbances.
  • The significant improvement by more nuanced and deeper linear and nonlinear decomposition provides insight for designing more effective and robust forecasting models.
@@ 2 Related Work
Time series forecasting is a critical area of research that finds applications in both industry and academia. With the powerful representation capability of neural networks, deep forecasting models have undergone a rapid development (Wang et al., 2024b; Lim & Zohren, 2021; Torres et al., 2021). Recent research endeavors have focused on segmenting the sequence into a series of patches (Nie et al., 2023; Zhang & Yan, 2023), better modeling the relationships between variables (Ng et al., 2022; Chen et al., 2024), the dynamic changes within a sequence (Wu et al., 2023; Du et al., 2023), or both (Yu et al., 2024; Liu et al., 2024a). Some works strive for more efficient forecasting solutions (Lin et al., 2024; Xu et al., 2024). Other efforts aim to revitalize existing architectures, such as RNN (Lin et al., 2023), Transformer (Liu et al., 2024b), TCN (donghao & wang xue, 2024), with new ideas, or to explore the potential of outstanding architectures from other domains, such as MLP-Mixer (Chen et al., 2023; Wang et al., 2024a), Mamba (Ahamed & Cheng, 2024; Wang et al., 2024c), Graph Neural Network (Yi et al., 2023; Shao et al., 2022), even Large Language Models (Jin et al., 2024; Liu et al., 2024c; Pan et al., 2024; Bian et al., 2024; Gruver et al., 2024), for application in time series forecasting. Notably, some efforts also begin to ponder the role of self-attention in time series forecasting (Ilbert et al., 2024).
@@
Deep time series models that are dedicated to model nonlinear patterns such as non-stationarity (Liu et al., 2022b), time-variant and time-invariant features (Liu et al., 2023), and frequency bias (Piao et al., 2024) have delivered outstanding performance in various domains. In contrast, recent advances have proved the importance of attention to linear patterns in time series (Toner & Darlow, 2024). For instance, DLinear (Zeng et al., 2023) and RLinear (Li et al., 2023a) achieve results comparable to, or even surpassing, many intricately designed nonlinear models in certain scenarios, using only simple linear layers. FITS (Xu et al., 2024) even achieves SOTA performance with merely $10k$ parameters. This suggests that balanced consideration of both linear and nonlinear patterns can be crucial for enhancing predictive performance in time series forecasting, which is unexplored in the current simple seasonal and trend decomposition.
@@ 3.1 Preliminary
Given a multivariate time series $X\in\mathbb{R}^{B\times C\times T}$ with a length of $T$ time steps, time series forecasting tasks are designed to predict its future $F$ time steps $\hat{Y}\in\mathbb{R}^{B\times C\times F}$, where $B$ is $batch\ size$, $C$ is the number of variate or channel ($C=1$ in univariate case), and $T$ represents the look-back window. We aim to make $\hat{Y}$ closely close to the ground truth $Y\in\mathbb{R}^{B\times C\times F}$.
@@
where $\lim\limits_{i\to\infty}R^{N}_{i}=0$. The modeling error will be $\delta=X-(\hat{L}_{1}+\hat{N}_{1}+\cdots+\hat{L}_{n}+\hat{N}_{n})=(L_{1}-\hat{L}_{1})+(N_{1}-\hat{N}_{1})+(L_{2}-\hat{L}_{2})+(N_{2}-\hat{N}_{2})+\cdots+(L_{S}-\hat{L}_{S})+(N_{S}-\hat{N}_{S})+\varepsilon$. By alternating the extraction of linear and nonlinear features with subtracting the extracted features from the original input representation, we ensure previously extracted features do not affect the extraction of subsequent features, guaranteeing the independence of the resulting linear and nonlinear patterns. Moreover, $\lim\limits_{i\to\infty}R^{N}_{i}=0$ ensures that all valuable information in the sequence is fully extracted if the decomposition level is deep enough, preventing any loss of information. This design takes the existing shallow RRD to a deeper level, introducing the extraction of nonlinear patterns. Such a refined design ensures that the model can achieve more robust forecasting results. Notably, if set the RRD level to $1$, then we get $Trend=IMF(X)=\hat{L}_{1}$, and $Seasonal=X-Trend=R^{L}_{1}$, where $IMF$ can be MOV, LD, or ESF, which is equivalent to the former seasonal and trend decomposition.
@@ 3.2 LiNo Pipeline
The structure of the LiNo framework is illustrated in Figure 2. Initially, we extract the whole series embedding, which is then processed through $N$ LiNo blocks to forecast future values of the time series. In this subsection, we will provide a detailed explanation of the whole series embedding, Li block, and No block, step by step.
@@
Following iTransformer (Liu et al., 2024b), we first map time series data $X\in\mathbb{R}^{B\times C\times T}$ from the original space to a new space to get $X_{embed}\in\mathbb{R}^{B\times C\times D}$ using a simple linear projection, where $D$ denotes the dimension of the layer. Such a design can better preserve the unique patterns of each variate.
@@
The fixed window size of the MOV and learnable 1D convolution kernel (LD), and the non-learnability of the exponential smoothing function (ESF), prevent them from fully extracting all linear patterns. So we introduce a learnable autoregressive model with a full receptive field to replace them, where MOV, LD, and ESF are its subset.
Given its input feature $H_{i}\in\mathbb{R}^{B\times C\times D}$ where $H_{0}=X_{embed}$, we get $i$-th linear pattern $L_{i}\in\mathbb{R}^{B\times C\times D}$ by:
$\begin{array}[]{c}\hat{L}_{i}[:,c,d]=\phi_{i}[c,1]*H_{i}[:,c,1]+\phi_{i}[c,2]*H_{i}[:,c,2]\\ +\cdots+\phi_{i}[c,d]*H_{i}[:,c,d]+\beta_{i}[c],\\ L_{i}=\textbf{Dropout}(\hat{L}_{i}).\end{array}$ (2)
Here, $\phi_{i}\in\mathbb{R}^{C\times D}$ represents the autoregressive coefficients, $\beta_{i}\in\mathbb{R}^{C}$ denotes the white noise error term. We assume the noise error term of each channel across all time stamps to be constant. So the whole process of extracting the linear part of the input feature can be easily deployed in a convolution fashion by setting the weight of the convolution kernel to $\phi_{i}$, and the weight of bias to $\beta_{i}$. Before applying the convolution, we pad the input feature $\mathbf{H}_{i}$. The padded input feature $\mathbf{H}_{i}^{\prime}$ becomes
$\mathbf{H}_{i}^{\prime}[:,:,t]=\begin{cases}\mathbf{H}_{i}[:,:,t-D],&\text{for }t\geq D\\ 0,&\text{for }t<0\end{cases}.$ (3)
We perform convolutions on each channel independently across the last dimension of $\mathbf{H}_{i}^{\prime}$. Following this, we apply dropout for generalization. The linear prediction $P^{li}_{i}\in\mathbb{R}^{B\times C\times F}$ for this level is obtained by mapping the extracted linear component $L_{i}$. Subsequently, we pass $R^{L}_{i}=H_{i}-L_{i}$ to the next stage for nonlinear pattern modeling.
@@
Typical nonlinear time series characteristics include temporal variations, frequency information, inter-series dependencies, etc. While all of these factors are crucial for precise forecasting, current nonlinear models can only address one (Wu et al., 2023; Nie et al., 2023; Zhang et al., 2022a; Piao et al., 2024) or two (Liu et al., 2024b; Li et al., 2023b; Chen et al., 2023) of these aspects. Therefore, we design a No block that can handle all these characteristics simultaneously.
Given a feature $R^{L}_{i}\in\mathbb{R}^{B\times C\times D}$, both temporal variation patterns and frequency information are accessible through linear projection. Hence, we start by applying a linear projection in the time domain to obtain temporal variation patterns $N^{T}_{i}\in\mathbb{R}^{B\times C\times D}$ in a classical manner. Correspondingly, we apply a linear projection in the frequency domain, which is transformed from $R^{L}_{i}$ using the Fast Fourier Transform (FFT), following FITS (Xu et al., 2024). Thereafter, the frequency domain representation is converted back using the Inverse Fast Fourier Transform (IFFT) to obtain frequency information patterns $N^{F}_{i}\in\mathbb{R}^{B\times C\times D}$. To leverage the complementary strengths of both domains, features extracted from both the time and frequency domains are fused and activated by: $N^{TF}_{i}=\textbf{Tanh}(N^{T}_{i}+N^{F}_{i})$.
To model inter-series dependencies, we first normalize $N^{TF}_{i}$ across the channel dimension using a softmax function. We then compute a weighted mean by multiplying the softmax weights with $N^{TF}_{i}$ and summing over the channel dimension. This weighted mean is repeated to match and concatenate with $R^{L}_{i}$. The concatenated result is passed through a Feedforward Network to obtain inter-series dependencies information $N^{C}_{i}$.
To integrate temporal variations, frequency information, and inter-series dependencies, we first apply Layer Normalization to the sum of $N^{TF}_{i}$ and $N^{C}_{i}$, resulting in $N^{TFC}_{i}$. A subsequent Feedforward Network is applied, and the result is added back to $N^{TFC}_{i}$, followed by a final Layer Normalization to produce the overall nonlinear pattern $N_{i}$. The nonlinear prediction $P^{no}_{i}\in\mathbb{R}^{B\times C\times F}$ for this level is then obtained by mapping the extracted nonlinear part $N_{i}$. Finally, $R^{N}_{i}=R^{L}_{i}-N_{i}$ is passed to the next LiNo block, where $H_{i+1}=R^{N}_{i}$.
@@
We used ReVIN (Kim et al., 2022) to counter the distribution problem following iTransformerLiu et al. (2024b). The input first performs an Instance Normalization (Ulyanov, 2016) before being embedded. The final output is reversed using the Mean and Standard Deviation of Instance Normalization. The final prediction result is aggregated from multi-level by:

### v4

@@ Abstract
Forecasting models are pivotal in a data-driven world with vast volumes of time series data that appear as a compound of vast Linear and Nonlinear patterns. Recent deep time series forecasting models struggle to utilize seasonal and trend decomposition to separate the entangled components. Such a strategy only explicitly extracts simple linear patterns like trends, leaving the other linear modes and vast unexplored nonlinear patterns to the residual. Their flawed linear and nonlinear feature extraction models and shallow-level decomposition limit their adaptation to the diverse patterns present in real-world scenarios. Given this, we innovate Recursive Residual Decomposition by introducing explicit extraction of both linear and nonlinear patterns. This deeper-level decomposition framework, which is named LiNo, captures linear patterns using a Li block which can be a moving average kernel, and models nonlinear patterns using a No block which can be a Transformer encoder. The extraction of these two patterns is performed alternatively and recursively. To achieve the full potential of LiNo, we develop the current simple linear pattern extractor to a general learnable autoregressive model, and design a novel No block that can handle all essential nonlinear patterns. Remarkably, the proposed LiNo achieves state-of-the-art on thirteen real-world benchmarks under univariate and multivariate forecasting scenarios. Experiments show that current forecasting models can deliver more robust and precise results through this advanced Recursive Residual Decomposition. Code is available at this Repository: https://github.com/Levi-Ackman/LiNo.
Time Series Forecasting, Linear, Nonlinear Time series
@@ 1 Introduction
Time series forecasting (TSF) is a long-established task (Lim & Zohren, 2021; Wang et al., 2024b), with a wide range of applications (Zhou et al., 2022a; Liu et al., 2022a; Piao et al., 2024). Notably, numerous deep learning methods have been employed to address the TSF problem, utilizing architectures such as Recurrent Neural Networks (RNNs) (Elman, 1990; Lin et al., 2023), Temporal Convolutional Networks (TCNs) (Donghao & Xue, 2024; Wu et al., 2023), Multilayer Perceptron (MLP) (Liu et al., 2023; Challu et al., 2023) and Transformers (Liu et al., 2022b; Kitaev et al., 2020). Recent advancements in the time series forecasting community have suggested that seasonal (nonlinear) and trend (linear) decomposition can enhance forecasting performance (Zhang et al., 2022b; Wu et al., 2021). This is typically performed once, with the trend component being extracted using methods such as the simple moving average kernel (MOV) (Zhou et al., 2022b; Wang et al., 2023), the exponential smoothing function (ESF) (Woo et al., 2022), or a learnable 1D convolution kernel (LD) (Yu et al., 2024). The seasonal component is then obtained by subtracting the trend part from the original time series.
@@
Figure 1: Example of the multi-level linear and nonlinear patterns in real-world time series. We take the ETTh2 dataset as an example and decompose a raw time series (Raw) into four signals through linear and nonlinear patterns decomposition. Linear 1&2 is the obtained linear patterns using the proposed Li block, and Nonlinear 1&2 is the obtained nonlinear patterns using No block. In other words, the raw time series (red) is the sum of the four signals below.
@@
Real-world time series data often exhibit more complex structures, combining multiple levels of linear and nonlinear characteristics (Stock & Watson, 1998; Dama & Sinoquet, 2021; Agrawal & Adhikari, 2013). This suggests that they are the result of the additive combination of various linear and nonlinear components, as illustrated in Figure 1. Simple seasonal and trend decomposition faces three main challenges when deployed in real-world time series data. Firstly, these methods rely on simple techniques (MOV, LD, ESF). MOV and LD employed a fixed window size, which cannot fully extract all linear patterns, while the non-learnability of ESF limits its performance. They can only extract simple linear features such as trends, while other linear features like cyclic patterns and autoregression (Chatfield & Xing, 2019) remain underutilized. Secondly, we argue the obtained seasonal part is actually a residual consisting of all the unextracted linear and nonlinear information. Without further separating the nonlinear part from the residual severely hinders its extraction. Since it’s extremely challenging for deep models to extract useful information from such mixtures (Bengio et al., 2013; Tishby & Zaslavsky, 2015). Another problem is the design of current nonlinear models, which mainly focus on one or two types of nonlinear patterns (temporal variations, frequency information, inter-series dependencies, etc). Such designs cannot satisfy our requirements in the real world. Thirdly, their shallow-level decomposition is incompatible with the multi-level characteristics of real-world time series. In contrast, numerous studies have shown that deeper-level decomposition can lead to better time series analysis (Huang et al., 1998; Rilling et al., 2003). Therefore, it is crucial to develop advanced decomposition methods capable of multi-granularity separation of various modes, and better linear and nonlinear pattern extractors to provide a more nuanced understanding of the signal’s structure and potentially improve forecasting performance.
@@
To address the challenges above, we propose adopting Recursive Residual Decomposition (RRD), a method used in Empirical Mode Decomposition (EMD) (Huang et al., 1998; Rilling et al., 2003), to decompose a time series into multiple patterns. This process is performed recursively. Each pattern is extracted based on the residual obtained by subtracting previously extracted patterns from the original signal, utilizing Intrinsic Mode Functions (IMF) to identify similar characteristics. Specifically, We can employ a Li block (MOV, LD, ESF, etc.) as the IMF for extracting linear patterns and a No block (Transformer, RNN, etc.) as the IMF for extracting nonlinear patterns. Then, the RRD is performed on a deeper level. We denote the proposed overall framework as LiNo. To fully realize RRD’s potential, we propose an advanced LiNo, Specifically, we enhance the Li block to a general learnable autoregressive model with a full receptive field and propose a novel No block capable of modeling essential nonlinear features, such as temporal variations, period information, and inter-series dependencies in multivariate forecasting.
In summary, our contributions can be delineated as follows:
@@
  • Proposed No block demonstrates better nonlinear pattern extraction ability than current SOTA nonlinear models, such as iTransformer and TSMixer.
  • LiNo consistently delivers top-tier performance in multivariate and univariate forecasting scenarios, demonstrating robust resilience against noise disturbances.
@@ 2 Related Work
Time series forecasting is a critical area of research that finds applications in both industry and academia. With the powerful representation capability of neural networks, deep forecasting models have undergone a rapid development (Wang et al., 2024b; Lim & Zohren, 2021). Recent research endeavors have focused on segmenting the sequence into a series of patches (Nie et al., 2023; Zhang & Yan, 2023), better modeling the relationships between variables (Ng et al., 2022; Chen et al., 2024), the dynamic changes within a sequence (Wu et al., 2023; Du et al., 2023), or both (Yu et al., 2024; Liu et al., 2024a). Some works strive for more efficient forecasting solutions (Lin et al., 2024; Xu et al., 2024). Other efforts aim to revitalize existing architectures, such as RNN (Lin et al., 2023), Transformer (Liu et al., 2024b), TCN (Donghao & Xue, 2024), with new ideas, or to explore the potential of outstanding architectures from other domains, such as MLP-Mixer (Chen et al., 2023; Wang et al., 2024a), Mamba (Ahamed & Cheng, 2024; Wang et al., 2024c), Graph Neural Network (Yi et al., 2023; Shao et al., 2022), even Large Language Models (Jin et al., 2024; Liu et al., 2024c), for application in time series forecasting.
@@
Deep time series models that are dedicated to model nonlinear patterns such as non-stationarity (Liu et al., 2022c), time-variant and time-invariant features (Liu et al., 2023), and frequency bias (Piao et al., 2024) have delivered outstanding performance in various domains. In contrast, recent advances have proved the importance of attention to linear patterns in time series (Toner & Darlow, 2024). For instance, DLinear (Zeng et al., 2023) and RLinear (Li et al., 2023a) achieve results comparable to, or even surpassing, many intricately designed nonlinear models in certain scenarios, using only simple linear layers. FITS (Xu et al., 2024) even achieves SOTA performance with merely $10k$ parameters. This suggests that balanced consideration of both linear and nonlinear patterns can be crucial for enhancing predictive performance in time series forecasting, which is unexplored in the current simple seasonal and trend decomposition.
Decomposition based on Residual.
Autoformer (Wu et al., 2021) first explored a rudimentary seasonal (Nonlinear) and trend (Linear) decomposition based on residual. The trend part is extracted using linear models like a moving average kernel, and the seasonal part is the residual of subtracting the input feature using the trend part. The earlier and more pioneering residual-based decomposition in deep learning time series was explored in N-BEATS (Oreshkin et al., 2020). N-BEATS is built using stacked residual blocks, where each block refines the forecast by predicting based on the residual errors of previous predictions. Instead of directly modeling the time series components (e.g., trend or seasonality), the model iteratively adjusts the residuals at each step.
Our LiNo develops the decomposition strategy in Autoformer by explicitly capturing both seasonal (Nonlinear) and trend (Linear) components and performing at a deeper level. Though LiNo and N-BEATS share some similarities, such as both employing the concept of Recursive Residual Decomposition (RRD), they are fundamentally different. First, N-BEATS does not explicitly capture linear and nonlinear separately. Second, N-BEATS cannot handle more complex multivariate time series. Third, the role of RRD differs in each model: while N-BEATS uses RRD to refine the final prediction based on residual errors from previous predictions, LiNo employs RRD to capture more nuanced linear and nonlinear patterns. To empirically compare their differences, we designed a model based on the N-BEATS framework, where the internal components are identical to those of the proposed LiNo. In the subsequent experiments, we will denote this model as Mu. A more detailed comparison and analysis of LiNo and N-BEATS is left in Appendix D.
@@ 3.1 Preliminary
Given a multivariate time series $X\in\mathbb{R}^{C\times T}$ with a length of $T$ time steps, time series forecasting tasks are designed to predict its future $F$ time steps $\hat{Y}\in\mathbb{R}^{C\times F}$, where $C$ is the number of variate or channel ($C=1$ in univariate case), and $T$ represents the look-back window. We aim to make $\hat{Y}$ close to the ground truth $Y\in\mathbb{R}^{C\times F}$.
@@
where $\lim\limits_{i\to\infty}R^{N}_{i}=0$. The modeling error will be $\delta=X-(\hat{L}_{1}+\hat{N}_{1}+\cdots+\hat{L}_{n}+\hat{N}_{n})=(L_{1}-\hat{L}_{1})+(N_{1}-\hat{N}_{1})+(L_{2}-\hat{L}_{2})+(N_{2}-\hat{N}_{2})+\cdots+(L_{S}-\hat{L}_{S})+(N_{S}-\hat{N}_{S})+\varepsilon$. By alternating the extraction of linear and nonlinear features with subtracting the extracted features from the original input representation, we ensure previously extracted features do not affect the extraction of subsequent features, guaranteeing the independence of the resulting linear and nonlinear patterns. Moreover, $\lim\limits_{i\to\infty}R^{N}_{i}=0$ ensures that all valuable information in the sequence is fully extracted if the decomposition level is deep enough, preventing any loss of information. This design takes the existing shallow RRD to a deeper level, introducing the extraction of nonlinear patterns. Such a refined design ensures the model achieves more robust forecasting results. Notably, if set the RRD level to $1$, then we get $Trend=IMF(X)=\hat{L}_{1}$, and $Seasonal=X-Trend=R^{L}_{1}$, where $IMF$ can be MOV, LD, or ESF, which is equivalent to the former seasonal and trend decomposition.
@@ 3.2 LiNo Pipeline
The structure of the LiNo framework is illustrated in Figure 2 (Middle). Initially, we extract the whole series embedding, which is then processed through $N$ LiNo blocks to forecast future values of the time series. In this subsection, we will explain the whole series embedding, Li block, and No block, step by step.
@@
Following iTransformer (Liu et al., 2024b), we first map time series data $X\in\mathbb{R}^{C\times T}$ from the original space to a new space to get $X_{embed}\in\mathbb{R}^{C\times D}$ using a simple linear projection, where $D$ denotes the dimension of the layer. Such a design can better preserve the unique patterns of each variate.
@@
The fixed window size of the MOV and learnable 1D convolution kernel (LD), and the non-learnability of the exponential smoothing function (ESF), prevent them from fully extracting all linear patterns. So we introduce a learnable autoregressive model (AR) with a full receptive field to replace them, where MOV, LD, and ESF are its subsets.
The structure of Li Block is shown in Figure 2 (right). Given its input feature $H_{i}\in\mathbb{R}^{C\times D}$ where $H_{0}=X_{embed}$, we get $i$-th linear pattern $L_{i}\in\mathbb{R}^{C\times D}$ by:
$\begin{array}[]{c}\hat{L}_{i}[c,d]=\phi_{i}[c,1]*H_{i}[c,1]+\phi_{i}[c,2]*H_{i}[c,2]\\ +\cdots+\phi_{i}[c,d]*H_{i}[c,d]+\beta_{i}[c],\\ L_{i}=\textbf{Dropout}(\hat{L}_{i}).\end{array}$ (2)
Here, $\phi_{i}\in\mathbb{R}^{C\times D}$ represents the autoregressive coefficients, $\beta_{i}\in\mathbb{R}^{C}$ denotes bias term. The whole process of extracting the linear part of the input feature can be easily deployed in a convolution fashion by setting the weight of the convolution kernel to $\phi_{i}$, and the weight of bias to $\beta_{i}$. Before applying the convolution (AR model), we pad the input feature $\mathbf{H}_{i}$ before convolution to make sure $H_{i}$ and $L_{i}$ have the same scale. The padded input feature $\hat{H}_{i}$ becomes,
$\hat{H}_{i}[:,t]=\begin{cases}{H}_{i}[:,t-D],&\text{for }t\geq D\\ 0,&\text{for }t<D\end{cases}.$ (3)
We perform convolutions on each channel independently across the last dimension of $\hat{H}_{i}$. Following this, we apply dropout for generalization. The linear prediction $P^{li}_{i}\in\mathbb{R}^{C\times F}$ for this level is obtained by mapping the extracted linear component $L_{i}$. Subsequently, we pass $R^{L}_{i}=H_{i}-L_{i}$ to the next stage for nonlinear pattern modeling.
@@
Typical nonlinear time series characteristics include temporal variations, frequency information, inter-series dependencies, etc. While all of these factors are crucial for precise forecasting, current nonlinear models can only address one (Wu et al., 2023; Nie et al., 2023; Zhang et al., 2022a; Piao et al., 2024) or two (Liu et al., 2024b; Li et al., 2023b; Chen et al., 2023) of these aspects. Therefore, we designed a No block that simultaneously handles all these characteristics.
As in Figure 2 (left), given a feature $R^{L}_{i}\in\mathbb{R}^{C\times D}$, both temporal variation patterns and frequency information are accessible through linear projection. Hence, we start by applying a linear projection in the time domain to obtain temporal variation patterns $N^{T}_{i}\in\mathbb{R}^{C\times D}$ in a classical manner. Correspondingly, we apply a linear projection in the frequency domain, which is transformed from $R^{L}_{i}$ using the Fast Fourier Transform (FFT), following FITS (Xu et al., 2024). After that, the frequency domain representation is converted back, using the Inverse Fast Fourier Transform (IFFT) to obtain frequency information patterns $N^{F}_{i}\in\mathbb{R}^{C\times D}$. To leverage the complementary strengths of both domains, features extracted from both the time and frequency domains are fused and activated by: $N^{TF}_{i}=\textbf{Tanh}(N^{T}_{i}+N^{F}_{i})$.
To model inter-series dependencies, we first normalize $N^{TF}_{i}$ across the channel dimension using a softmax function, as in Figure 2 (upper left). We then compute a weighted mean by multiplying the softmax weights with $N^{TF}_{i}$ and summing over the channel dimension. This weighted mean is repeated to match and concatenate with $N^{TF}_{i}$. The concatenated result $\hat{N}^{C}_{i}$ is passed through an MLP to obtain inter-series dependencies information $N^{C}_{i}$. We note that a recent work, SOFTS (Han et al., 2024), has adopted a similar approach to ours for modeling inter-series dependencies. A detailed comparative analysis is provided in Appendix E.
To integrate temporal variations, frequency information, and inter-series dependencies, we first apply Layer Normalization to the sum of $N^{TF}_{i}$ and $N^{C}_{i}$, resulting in $N^{TFC}_{i}$. A subsequent MLP is applied, and the result is added back to $N^{TFC}_{i}$, followed by a final Layer Normalization to produce the overall nonlinear pattern $N_{i}$. The nonlinear prediction $P^{no}_{i}\in\mathbb{R}^{C\times F}$ for this level is then obtained by mapping the extracted nonlinear part $N_{i}$. Finally, $R^{N}_{i}=R^{L}_{i}-N_{i}$ is passed to the next LiNo block, where $H_{i+1}=R^{N}_{i}$.
@@
We used RevIN (Kim et al., 2022) to counter the distribution problem following iTransformer(Liu et al., 2024b). The final prediction result is aggregated from multi-level by: